OpenStudy (anonymous):

Condense ln(x)-4[ln(x+2)+ln(x-2)]

hero (hero):

\(\ln(a) + \ln(b) = \ln(ab)\)

hero (hero):

\(\ln(a) - \ln(b) = \ln\left(\frac{a}{b}\right)\)